### v1

# VL-Nav: Real-time Vision-Language Navigation with Spatial Reasoning

Abstract
Vision-language navigation in unknown environments is crucial for mobile robots. In scenarios such as household assistance and rescue, mobile robots need to understand a human command, such as “find a person wearing black”. We present a novel vision-language navigation (VL-Nav) system that integrates efficient spatial reasoning on low-power robots. Unlike prior methods that rely on a single image-level feature similarity to guide a robot, we introduce the heuristic-vision-language (HVL) spatial reasoning for goal point selection. It combines pixel-wise vision-language features and heuristic exploration to enable efficient navigation to human-instructed instances in various environments robustly. We deploy VL-Nav on a four-wheel mobile robot and conduct comprehensive navigation tasks in various environments of different scales and semantic complexities, indoors and outdoors. Remarkably, VL-Nav operates at a real-time frequency of 30 Hz with a Jetson Orin NX, highlighting its ability to conduct efficient vision-language navigation. Experimental results show that VL-Nav achieves an overall success rate of 86.3%, outperforming previous methods by 44.15%.
Figure 1: We propose VL-Nav, a real-time zero-shot vision-language navigation approach with spatial reasoning that integrates pixel-wise vision-language features and heuristic cues for mobile robots. (a) Hallway: The wheeled robot is tasked with “find a man in gray” in a hallway. Unlike the classical frontier-based method (red line) and VLFM (green line), VL-Nav (blue line) leverages pixel-wise vision-language (VL) features from the “gray cloth” cue for spatial reasoning, selecting the nearby frontier point as the goal and successfully locating the missing person. The value map shows that the “gray cloth” VL cue prioritizes the right-side area, marked by yellow square points. (b) Apartment: The robot is tasked with “Go to the tall white trash bin.” It detected two different-sized white trash bins in bottom camera observation. However, it assigns a higher confidence score (0.98) to the taller bin than the shorter one (0.48). These pixel-wise VL features are incorporated into the spatial distribution to select the correct goal point, guiding the robot toward the taller bin.
I Introduction
Efficient navigation following human instructions in unseen environments is critical for autonomous robots, with various applications spanning from home assistance [28, 31, 32] to planetary exploration [25, 35, 27]. In such scenarios, robots are tasked with exploring and identifying targets instructed by humans, a challenge we define as vision-language navigation (VLN). Suppose a man wearing gray is reported missing and a command “find a man in gray” is given, the robot should not engage in random exploration. Instead, it should prioritize directions with visual features showing a stronger correlation to the language cues, such as detecting “gray cloth” within its field of view. Real-world VLN requires a navigation system that can: (1) interpret pixel-wise vision-language features, (2) adapt and perform robustly across various environments, and (3) operate in real-time on low-power platforms.
However, there are still no VLN systems that can fully address the three capabilities. Existing approaches can be broadly categorized into classical, end-to-end learning, and modular learning methods [12]. Classical approaches [9, 10, 29, 33], while efficient, struggle to integrate vision-language features. End-to-end learning methods [17, 24, 7, 26, 22] show promise but are computationally expensive, prone to overfitting in simulation, and exhibit poor generalization to out-of-distribution scenarios. Modular learning approaches [3, 21, 13, 5, 4], on the other hand, demonstrate robust real-world performance [12]. However, they often depend on extensive real-world robot training data and lack the capacity for human-like reasoning. The advent of vision-language models (VLMs) and large language models (LLMs) has further enhanced modular navigation approaches. For instance, Vision-Language Frontier Maps (VLFM) [36] utilized VLMs to extract language-grounded features directly from RGB images. This enables the creation of a semantic map, guiding exploration based on human semantic knowledge. However, their reliance on computationally intensive models limits their deployability on low-power platforms. Moreover, VLFM relies heavily on a single image-level feature similarity for goal selection, limiting its ability to exploit fine-grained vision-language cues.
To bridge this gap, we present vision-language navigation (VL-Nav), a novel navigation framework optimized for low-power robots, achieving zero-shot VLN at 30 Hz with an onboard computer. VL-Nav employs heuristic vision-language (HVL) spatial reasoning on both frontier-based and instance-based target points. VL-Nav initially generates frontier-based target points from the dynamic occupancy map using partial frontier detection. It confines the search to a manageable field of view, thereby reducing computational overhead. Additionally, instance-based target points are incorporated to emulate human search patterns, enabling the robot to approach and verify potential target objects, thereby enhancing the success rate. To select the most informative goal point, we introduce HVL spatial reasoning. This technique first transforms pixel-wise vision-language features into a spatial scoring distribution using a Gaussian mixture model. Each target point is then assigned a vision-language semantic score based on this distribution. Subsequently, a curiosity-driven heuristic weighting, which encourages the robot to explore unknown areas, is applied to adjust these scores into HVL scores. The target point with the highest HVL score is ultimately selected as the goal point. This HVL spatial reasoning process ensures that the selected goal point not only closely aligns with the human description but also guides the robot to explore unknown areas.
Once a goal is selected, VL-Nav utilizes a classic planner for real-time obstacle-aware path planning, enabling seamless adaptation to partially known environments. By combining pixel-wise vision-language features with heuristics via the novel HVL spatial reasoning, VL-Nav surpasses all the baselines, achieving human-like intelligent navigation while remaining computationally feasible for real-world deployment.
Our main contributions are as follows:
  • We propose VL-Nav, an efficient VLN system optimized for low-power robots, achieving robust real-time performance at 30 Hz with an onboard computer.
  • We introduce the heuristic-vision-language (HVL) spatial reasoning combining pixel-wise vision-language features and heuristic exploration for efficient and robust VLN..
  • We conduct comprehensive evaluations across four real-world settings, showing that VL-Nav outperforms prior methods by 44.15% in diverse environments.
II Related Works
Robot navigation methods can be broadly categorized into classical, end-to-end learning, and modular learning methods.
Classical Approaches:
Classical methods for navigation, including Simultaneous Localization and Mapping (SLAM)-based techniques, have been extensively studied over the last three decades [9, 10, 29, 33]. These methods typically build geometric maps of an environment using depth sensors or monocular RGB cameras [30, 19, 8, 23], while simultaneously localizing the robot relative to the growing map. Exploration is often guided by heuristic strategies like frontier-based methods [33, 11], while analytical planners are employed for obstacle avoidance and path planning. Although effective for traditional navigation tasks, these classical methods lack the ability to integrate vision-language features, making them insufficient for human-collaboration tasks like vision-language navigation.
End-to-End Learning:
End-to-end learning methods directly map sensory inputs to navigation actions using deep neural networks, typically trained with imitation learning (IL) or reinforcement learning (RL) losses [17, 24, 7, 26, 22]. Despite their ability to learn semantic priors for goal-directed exploration, end-to-end approaches suffer from significant limitations. These include high computational cost, reliance on large-scale training data, and a lack of interpretability. Furthermore, these methods often struggle to generalize beyond simulation due to domain gaps, as evidenced by significant performance drops when transitioning to real world [12].
Figure 2: An overview of the VL-Nav pipeline. VL-Nav processes inputs including prompts, RGB images, odometry poses, and LiDAR scans. The Vision-Language (VL) module conducts open-vocabulary pixel-wise detection to identify areas and objects related to the prompt, generating instance-based target points. Concurrently, the map module performs terrain analysis and manages a dynamic occupancy map. Frontier-based target points are then identified based on this occupancy map, along with the instance points, forming a candidate points pool. VL-Nav employs heuristic-vision-language (HVL) spatial reasoning to select the most effective goal point from this pool for path planning.
Modular Learning:
Modular learning approaches seek to combine the advantages of classical and end-to-end methods by replacing specific components of the classical pipeline with learned modules. This modularity allows for the integration of semantic priors while maintaining system interpretability and efficiency [16, 18, 2, 1]. For example, semantic maps generated from RGB-D data have been used to guide navigation toward specific objects [3, 20, 21, 13]. Modular methods have shown promise in addressing Sim-to-Real transfer challenges by abstracting raw sensor data into higher-level representations, mitigating the impact of domain gaps [18, 1, 2]. However, existing modular approaches often rely heavily on real-world training data and struggle to utilize vision-language features.
Foundation Models for Navigation:
Recent advances in foundation models like VLMs and LLMs have further enhanced semantic navigation by incorporating natural language understanding into modular learning approaches. The Vision-Language Frontier Maps (VLFM) method [36] exemplifies this trend by leveraging a pre-trained VLM to extract semantic values directly from RGB images, enabling zero-shot semantic prior understanding for frontier-based exploration. VLFM achieves SOTA performance in the simulated environment, addressing the limitations of prior frontier-based and goal-oriented modular methods [33, 11, 3, 15]. However, VLFM’s reliance on computationally intensive VLMs restricts its deployability on resource-constrained platforms, and its focus on a single semantic similarity score for each observation does not adequately account for pixel-wise vision-language understanding and spatial reasoning, which is crucial for VLN. Their real-robot demonstration involves mounting a high-power-consumption RTX 4090 gaming laptop on the quadrupedal robot to handle computation-intensive modules, which is an impractical design for most real-world applications.
Strengths of VL-Nav:
VL-Nav significantly advances vision-language navigation (VLN) by addressing the limitations of classical, end-to-end, and modular learning methods. It employs a rolling occupancy grid and partial frontier detection, reducing computational overhead and enhancing adaptability in dynamic environments on low-power platforms. The integration of instance-based target points with heuristic vision-language (HVL) spatial reasoning enables pixel-wise vision-language feature interpretation and context-aware goal selection. Unlike systems reliant on heavy computational models, VL-Nav optimizes for resource-constrained platforms and maintains real-time performance at 30 Hz. This makes VL-Nav not only practical for real-world applications but also superior in navigating complex environments efficiently.
III Methodology
Figure 3: A brief illustration of VL Scoring. The pixel-wise open-vocabulary detection results are transferred into the spatial distibution via the Gaussian mixture model regularized by the FOV weighting (the gray arrows in the figure). Then the frontier-based and the instance-based target points will be assigned with VL score based on the distribution Equation (3).
This section presents our VL-Nav pipeline (Fig. 2), which leverages a rolling occupancy grid map, frontier-based and instance-based target points, and Heuristic-Vision-Language (HVL) spatial reasoning to select and publish a navigation goal. This goal is not only aligned with the given prompt but also leads the robot to explore unknown areas more effectively.
III-A Rolling Occupancy Map: Occupancy Update
We represent the environment as a 2D occupancy grid $\mathcal{G}$, where each cell may be free $(0)$, unknown $(-1)$, or occupied $(100)$. Upon receiving new sensor data (a merged obstacle-and-terrain cloud $\mathcal{P}$), we update $\mathcal{G}$ using the following steps:
  1. 1.
    Expand if needed. If any newly observed points lie outside the current map boundaries, expand the grid in place to preserve historical data.
  2. 2.
    Clear stale obstacles in forward FOV. For each cell marked occupied within the robot’s forward field of view (FOV), check whether it still corresponds to a point in $\mathcal{P}$. Any “stale” obstacle cells are re-labeled as free.
  3. 3.
    Inflate new obstacles. For each point in $\mathcal{P}$ within the sensor range $R$, mark the corresponding cell as occupied, along with a local neighborhood determined by an inflation radius $r_{\mathrm{inflation}}$.
  4. 4.
    Raycast unknown cells to free. Perform line-of-sight raycasting from the robot’s pose $(x_{r},y_{r},\theta_{r})$ across its angular range $\theta\in[\theta_{r}-\tfrac{\mathrm{hfov}}{2},\;\theta_{r}+\tfrac{\mathrm{hfov}}{2}]$. Any unknown cells $(-1)$ intersected by a ray are marked free $(0)$.
Algorithm 1 summarizes these main steps, returning the updated grid $\mathcal{G}$. This occupancy update leverages consistent line-of-sight logic and inflation to support safe navigation.
Algorithm 1 Rolling Occupancy Map Update (Key Steps)
0:  Grid $\mathcal{G}$, merged point cloud $\mathcal{P}$, robot pose $(x_{r},y_{r},\theta_{r})$, inflation radius $r_{\mathrm{inflation}}$, half-FOV $\mathrm{hfov}/2$, sensor range $R$.
1:  $\textsc{ExpandIfNeeded}(\mathcal{G},\mathcal{P})$
[…]
5:  return  Updated grid $\mathcal{G}$
III-B Frontier-based Target Points
A cell $(m_{x},m_{y})$ is a frontier if it is free, $\mathcal{G}(m_{x},m_{y})=0$, and at least one neighbor is unknown, $\mathcal{G}(m_{x}+\Delta x,m_{y}+\Delta y)=-1$ for some $(\Delta x,\Delta y)$. After every grid update, only cells in the forward wedge are tested:
$\begin{split}\Bigl{|}\mathrm{Angle}\bigl{(}(m_{x},m_{y}),(x_{r},y_{r})\bigr{)}-\theta_{r}\Bigr{|}\;&\leq\;\frac{\mathrm{hfov}}{2},\\ \text{and}\quad\|(m_{x},m_{y})-(x_{r},y_{r})\|\;&\leq\;R.\end{split}$
A breadth-first search (BFS) clusters these frontier cells. Each cluster is represented by either a single centroid (for small clusters) or multiple sampled points (for large clusters) as the frontier-based target points.
III-C Instance-Based Target Points (IBTP)
A vision-language detector periodically reports candidate instance centers in the form $(q_{x},\;q_{y},\;\text{confidence})$, where $q_{x},q_{y}$ denote the estimated global coordinates for a potential target instance, and confidence quantifies how likely this detection is to match the desired instance. If confidence is larger than the detection threshold $\text{confidence is larger than the detection threshold }\tau_{\mathrm{det}}$, the candidate is retained; otherwise, it is discarded as too uncertain. Retained points will be downsampled via a voxel-grid filter if many candidates lie in close proximity.
This instance-based approach mimics human behavior when searching for an object: upon glimpsing something that might match the target, one naturally moves closer to confirm. VL-Nav likewise does not ignore intermediate detections. Instead, any candidate above a confidence threshold is treated as a valid goal candidate, allowing the robot to approach and verify whether it truly is an instance of interest. If the detection turns out to be incorrect, the robot continues exploring via frontiers or other instance cues, producing an accurate and robust strategy for zero-shot navigation toward a target instance.
III-D Heuristic-Vision-Language (HVL) Spatial Reasoning
Once frontier centroids and instance-based targets are gathered, the system computes an HVL score for each candidate goal $\mathbf{g}$. Let $(x_{r},y_{r})$ be the robot position, $\theta_{r}$ the heading, and $d(\mathbf{x}_{r},\mathbf{g})$ the Euclidean distance. The angular offset is
$\Delta\theta\;=\;\mathrm{Angle}\bigl{(}\mathbf{g},(x_{r},y_{r})\bigr{)}\;-\;\theta_{r}.$
VL Score:
As shown in Fig. 3, a vision-language (VL) score $S_{\mathrm{VL}}(\mathbf{g})$ translates pixel-wise vision-language features into a Gaussian-mixture distribution over the robot’s horizontal field of view (FOV). Suppose the open-vocabulary detection model identifies $K$ likely directions or detection regions, each parameterized by $\bigl{(}\mu_{k},\sigma_{k},\alpha_{k}\bigr{)}$. Here, $\mu_{k}$ indicates the mean offset angle within the FOV, $\sigma_{k}$ encodes the detection’s angular uncertainty (in our implementation, a fixed constant of $0.1$ for all detections), and $\alpha_{k}$ is a confidence-based weight that captures how important that region is. By accumulating each region’s contribution in a mixture model, $S_{\mathrm{VL}}(\mathbf{g})$ naturally biases the robot toward goals in the direction of high-confidence detections that match the language prompt.
From a human perspective, this process resembles observing different areas that related to the target instance with varying levels of certainty. For instance, if someone is looking for a “red chair” in a large room, they might spot multiple red silhouettes at the right side of their vision, plus a clearer view of something red near the center. Although peripheral glimpses (“out of the corner of the eye”) often carry lower confidence, they are not ignored. Instead, we naturally merge these partial observations into a mental sense of “the chair is probably in one of these directions.” The system follows a similar approach, assigning each detection region a Gaussian shape around $\mu_{k}$, scaled by $\sigma_{k}$ and weighted by $\alpha_{k}$. Inspired by VLFM [36], we then multiply by $C(\Delta\theta)$, a “view confidence” term that downweights detections with large angular offsets from the central field of view. The VL score is computed as:
$S_{\mathrm{VL}}(\mathbf{g})\;=\;\sum_{k=1}^{K}\alpha_{k}\;\exp\!\Bigl{(}-\tfrac{1}{2}\Bigl{(}\tfrac{\Delta\theta-\mu_{k}}{\sigma_{k}}\Bigr{)}^{2}\Bigr{)}\;\cdot\;C\bigl{(}\Delta\theta\bigr{)},$
$C(\Delta\theta)=\cos^{2}\!\biggl{(}\frac{\Delta\theta}{(\theta_{\mathrm{fov}}/2)}\cdot\frac{\pi}{2}\biggr{)}.$
Heuristic Cues:
We add two heuristic terms to guide navigation toward larger unexplored areas and prevent the system from repeatedly selecting faraway goals that offer only a marginal increase in VL score—thus reducing unnecessary back-and-forth movement which is extremely important for the large-scale environments like the hallway and the outdoor environments in our experiments.
(1) Distance Weighting. We define
$S_{\mathrm{dist}}(\mathbf{g})\;=\;\frac{1}{1+d(\mathbf{x}_{r},\mathbf{g})},$
so that nearer goals receive slightly higher scores. This factor is especially important on a real robot, as shorter travel distances can significantly reduce energy consumption while preventing needless wandering. Although the distance term alone does not guarantee an optimal path, it helps prioritize goals that can be reached sooner.
(2) Unknown-Area Weighting. We further encourage curiosity-driven exploration by measuring how many unknown cells lie around $\mathbf{g}$. First, we define the ratio of unknown as:
[…]
where a local BFS radiates outward from $\mathbf{g}$, counting how many cells are unknown versus how many are reachable (i.e., not blocked by obstacles). Larger $\mathrm{ratio}(\mathbf{g})$ implies that moving to $\mathbf{g}$ may reveal significant unknown space and get more information. To translate this raw ratio into a normalized weighting score, we apply an exponential mapping:
[…]
where $k$ is an adjustable parameter that controls how rapidly the score increases from 0 toward 1.
Combined HVL Score:
The final scores of frontier-based goals combine these three components:
$S_{\mathrm{HVL}}(\mathbf{g})\;=\;\Bigl{(}w_{\mathrm{dist}}\;S_{\mathrm{dist}}(\mathbf{g})\;+\;w_{\mathrm{VL}}\;S_{\mathrm{VL}}(\mathbf{g})\;\cdot\;S_{\mathrm{unknown}}(\mathbf{g})\Bigr{)},$
where $w_{\mathrm{VL}}$ and $w_{\mathrm{dist}}$ are scalar weights. Note that the VL score is assigned when a point is initially detected, whereas the heuristic terms (distance and unknown-area weighting) are evaluated at goal selection time—when the current distance and unknown ratio can be calculated. For instance-based goals, these heuristic cues are omitted because their primary purpose is verification rather than exploration.
III-E Select & Publish Goal Point
After evaluating HVL scores for all instance-based and frontier goals, the system picks the highest-scoring candidate. Algorithm 2 sketches the selection process, prioritizing instances first, then frontiers.
Algorithm 2 Select and Publish Goal Point
0:  Instance points $\{\mathbf{o}_{1},\dots\}$ with $VL\_score$, frontier points $\{\mathbf{f}_{1},\dots\}$ with $HVL\_score$, robot position $\mathbf{r}$, and threshold $\delta_{\mathrm{reached}}$ for discarding points too close to $\mathbf{r}$.
1:  $\textit{bestScore}\leftarrow-\infty$
[…]
3:  for all instances $\mathbf{o}_{i}$ do
4:     if $\|\mathbf{o}_{i}-\mathbf{r}\|>\delta_{\mathrm{reached}}$ then
[…]
6:        if $s>\textit{bestScore}$ then
7:           $\textit{bestScore}\leftarrow s$
8:           $\textit{bestGoal}\leftarrow\mathbf{o}_{i}$
9:        end if
10:     end if
11:  end for
[…]
13:     for all frontiers $\mathbf{f}_{j}$ do
14:        if $\|\mathbf{f}_{j}-\mathbf{r}\|>\delta_{\mathrm{reached}}$ then
[…]
16:           if $s>\textit{bestScore}$ then
17:              $\textit{bestScore}\leftarrow s$
18:              $\textit{bestGoal}\leftarrow\mathbf{f}_{j}$
19:           end if
20:        end if
21:     end for
22:  end if
[…]
24:     Publish bestGoal to /selected_goal_point
25:  end if
26:  return
If no candidate exceeds the distance threshold $\delta_{\mathrm{reached}}$, the robot withholds any new goal. Once a goal is published, a global planner updates obstacle polygons and replans a safe route. A local planner refines each global waypoint into short-horizon commands for real-time obstacle avoidance, supporting incremental exploration in unknown environments.
III-F Path Planning with FAR Planner
Once an HVL goal is selected, we hand off point-goal path planning to the FAR Planner [34], which represents obstacles polygonally and updates a visibility graph in real time. This enables efficient re-planning in partially unknown environments, often outperforming search- or sampling-based methods. A local planner refines FAR Planner’s waypoints into short-horizon velocity commands, ensuring swift reactions to new obstacles and seamless integration with VL-Nav’s zero-shot vision-language objectives.
IV Experiments
Figure 4: Four different real-world experiment environments.
(a) Office
(b) Apartment
(c) Outdoor
(d) Hallway
Figure 5: Top-down view of the trajectories comparison on the value maps with the detection results across the four different environments.
TABLE I: Vision-language navigation performance in 4 unseen environments (SR and SPL).
Method SR (%) SPL
Hallway Office Apartment Outdoor Hallway Office Apartment Outdoor
Frontier Exploration 40.0 41.7 55.6 33.3 0.239 0.317 0.363 0.189
VLFM [36] 53.3 75.0 66.7 44.4 0.366 0.556 0.412 0.308
VL-Nav w/o IBTP 66.7 83.3 70.2 55.6 0.593 0.738 0.615 0.573
VL-Nav w/o Heuristics 73.3 86.3 66.7 55.6 0.612 0.743 0.631 0.498
VL-Nav 86.7 91.7 88.9 77.8 0.672 0.812 0.733 0.637
IV-A Experimental Setting
We evaluate our approach in real-robot experiments against five methods: (1) classical frontier-based exploration, (2) VLFM [36], (3) VLNav without instance-based target points, (4) VLNav without heuristic terms, and (5) the full VLNav configuration. Because the original VLFM relies on BLIP-2 [14], which is too computationally heavy for real-time edge deployment, we use the YOLO-World [6] model instead to generate per-observation similarity scores for VLFM. Each method is tested under the same conditions to ensure a fair comparison of performance.
Environments:
We consider four distinct environments (shown in Fig. 4), each with a specific combination of semantic complexity (High, Medium, or Low) and size (Big, Mid, or Small). Concretely, we use a Hallway (Medium & Big), an Office (High & Mid), an Outdoor area (Low & Big), and an Apartment (High & Small). In each environment, we evaluate five methods using three language prompts, yielding a diverse range of spatial layouts and semantic challenges. This setup provides a rigorous assessment of each method’s adaptability.
Language-Described Instance:
We define nine distinct, uncommon human-described instances to serve as target objects or persons during navigation. Examples include phrases such as “tall white trash bin,” “there seems to be a man in white,” “find a man in gray,” “there seems to be a black chair,” “tall white board,” and “there seems to be a fold chair.” The variety in these descriptions ensures that the robot must rely on vision-language understanding to accurately locate these targets.
Robots and Sensor Setup: All experiments are conducted using a four-wheel Rover equipped with a Livox Mid-360 LiDAR. The LiDAR is tilted by approximately 23 degrees to the front to achieve a ±$\pm 30$ degrees vertical FOV coverage closely aligned with the forward camera’s view. An Intel RealSense D455 RGB-D camera, tilted upward by 7 degrees to detect taller objects, provides visual observation, though its depth data are not used for positioning or mapping. LiDAR measurements are a primary source of mapping and localization due to their higher accuracy. The whole VL-Nav system runs on an NVIDIA Jetson Orin NX on-board computer.
IV-B Main Results
We validate the proposed VL-Nav system in real-robot experiments across four distinct environments (Hallway, Office, Apartment, and Outdoor), each featuring different semantic levels and sizes. Building on the motivation articulated in Section I, we focus on evaluating VL-Nav’s ability to (1) interpret fine-grained vision-language features and conduct robust VLN, (2) explore efficiently in unfamiliar spaces across various environments, and (3) run in real-time on resource-constrained platforms. Fig. 5 presents a top-down comparison of trajectories and detection results on the value map.
Figure 6: Plots of performance in different environments sizes and semantic comlexities.
Overall Performance:
As reported in Table I, our full VL-Nav consistently obtains the highest Success Rate (SR) and Success weighted by Path Length (SPL) across all four environments. In particular, VL-Nav outperforms classical exploration by a large margin, confirming the advantage of integrating HVL spatial reasoning with partial frontier-based search rather than relying solely on geometric exploration.
Effect of Instance-Based Target Points (IBTP):
We note a marked improvement when enabling IBTP: the variant without IBTP lags behind, particularly in complex domains like the Apartment and Office. As discussed in Section III, IBTP allows VL-Nav to pursue and verify tentative detections with confidence above a threshold, mirroring human search behavior. This pragmatic mechanism prevents ignoring possible matches to the target description and reduces overall travel distance to confirm or discard candidate objects.
Heuristics Contributions:
The Heuristic Score is also significant to VL-Nav’s performance. It merges two key components:
  • Distance Weighting: Preventing easily select very far way goals to reduce travel time and energy consumption which is extremely important for the efficiency (metrics SPL) in the large-size environments.
  • Unknown-Area Heuristic: Rewards navigation toward regions that yield more information.
Our ablations reveal that removing the distance-scoring element (VL-Nav w/o heuristics) degrades both SR and SPL, particularly in the more cluttered environments. Meanwhile, dropping the instance-based target points (IBTP) similarly lowers performance, reflecting how each piece of HVL addresses a complementary aspect of semantic navigation.
Comparison to VLFM:
Although the VLFM approach [36] harnesses vision-language similarity value, it lacks the pixel-wise vision-language features, instance-based target points verification mechanism, and HVL-based spatial reasoning. Consequently, VL-Nav surpasses VLFM in both SR and SPL by effectively combining the pixel-wise vision language features and the heuristic cues via the HVL spatial reasoning. These gains are especially pronounced in semantic complex (Apartment) and open-area (Outdoor) environments, underscoring how our HVL spatial reasoning enhance vision-language navigation in complex settings and scenarios.
Summary of Findings:
In conclusion, the experimental results confirm that VL-Nav delivers superior vision-language navigation across diverse, unseen real-world environments. By fusing frontier-based target points detection, instance-based target points, and the HVL spatial reasoning for goal selection, VL-Nav balances semantic awareness and exploration efficiency. The system’s robust performance, even in large or cluttered domains, highlights its potential as a practical solution for zero-shot vision-language navigation on low-power robots.
V Discussion
V-A Sizes of Environments vs. SR and SPL
Our experiments grouped environments by size into three categories: Small (Apartment), Mid (Office), and Big (averaging Hallway and Outdoor). As shown in Table I and the line plots in Fig. 6, classical Frontier Exploration suffers a marked performance decline in larger spaces, frequently exceeding the path length limit. Its success rate drops from about 55.6% in the Small environment to only 36.7% in the Big, and its SPL falls from 0.564 to 0.214. This illustrates how naive 360° frontier-based approaches grow increasingly inefficient in larger, more open environments.
By contrast, VL-Nav remains robust across all sizes. Although there is a slight decrease when moving from Mid (91.7% SR, 0.852 SPL) to Big (82.3% SR, 0.677 SPL), its overall performance remains substantially higher than the baselines. A similar trend appears for the ablated VL-Nav variants (w/o IBTP, w/o Heuristic), which still surpass Frontier Exploration but fall behind the full VL-Nav system.
These findings underscore two key factors. First, HVL spatial reasoning allows our method to leverage fine-grained vision-language cues (e.g., the “gray cloth” in Fig. 1) while also prioritizing areas with substantial unknown space, thus avoiding the extreme inefficiencies of frontier-based exploration faces in large, open environments. Second, instance-based target points (IBTP) enable more targeted navigation, sustaining high success rates even as map size grows.
V-B Semantic Complexities of Environments vs. SR and SPL
We also vary the environments by semantic complexity, classifying them as Low (Outdoor), Medium (Hallway), and High (averaging Office and Apartment). An interesting trend emerges: all methods show better performance in more semantically rich environments. The likely explanation is that structured indoor spaces (e.g., with walls or furniture) provide stronger cues for the detection and segmentation models used by all five methods, leading to more accurate navigation.
These results underscore that VL-Nav gains more pronounced benefits as the environment offers more semantic context. Rather than being overwhelmed by clutter or complexity, the system capitalizes on language-guided instance-based detections and heuristic exploration. The fusion of the HVL spatial reasoning and instance-based target points provides a scalable strategy to locate human-specified targets, even in highly complex indoor domains.
V-C Dynamic Environment Handling
Handling dynamic environments is crucial for robust navigation in real-world scenarios, where changes such as moving obstacles and varying terrain can impact robot performance. VL-Nav incorporates several strategies to dynamically adapt to environmental changes, ensuring reliability and efficiency.
To handle the dynamic environment, we maintain a dynamic occupancy grid map that integrates dynamic terrain analysis and path planning within the FAR planner [34]. This allows the system to continuously update its understanding of the environment and adjust navigation strategies accordingly. The dynamic grid map ensures that real-time changes, such as shifting obstacles and terrain conditions, are accurately reflected, providing a more robust foundation for dynamic vision-language navigation in the social environment.
V-D Real-Robot Efficiency Optimization
Although VL-Nav is designed as a general navigation framework, we introduce several optimizations to ensure real-world feasibility and high performance on mobile robots.
Visual-Language Module Variants:
The first optimization involves selecting an efficient model variant for the visual-language (or object-detection) pipeline. In our tests, we ran multiple versions of YOLO-World [6] detectors with different input dimensions and TensorRT precision modes. For instance, one configuration used a $256\times 320$ input size in a standard GPU runtime, another used TensorRT (FP32) at the same resolution, and a third tested TensorRT (FP16) at $480\times 640$. Empirically, the original $256\times 320$ configuration (run without TensorRT) achieved about $23\%$ higher success rates than higher-resolution FP16 modes. This suggests that moderate input resolutions can strike a more favorable balance between detection accuracy and real-time performance, which is crucial for robust and efficient continuous on-robot inference.
Rolling Occupancy Grid vs. Fixed Large Grid:
A second optimization lies in how the occupancy grid is maintained. Rather than keeping a large, fixed-size global grid (common in many simulation environments), we implement a rolling grid on the real robot, dynamically shifting or expanding it only as new sensor data demand. This rolling approach keeps memory usage proportional to the active exploration area, lowering both storage overhead and BFS/cluster computations for partial frontier detection. Moreover, since real-world robots typically explore incrementally, discarding far-off map regions has minimal impact on decision-making.
Additional System-Level Improvements:
Finally, we incorporate various practical measures to ensure robust navigation under real-world constraints. For instance, the robot’s local planner and velocity commands are monitored to detect “stuck” situations; if the robot remains motionless beyond a time threshold, the system discards its current goal and reselects a new one. Similarly, sensor frequency and map update rates are tuned so that occupancy-grid expansions do not saturate available CPU/GPU resources. These measures, alongside the rolling-grid and efficient detection models, help keep VL-Nav lightweight and reactive.
VI Limitation and Conclusion
VI-A Limitations
Despite the promising results demonstrated by VL-Nav, several limitations remain. One major challenge is the system’s ability to handle complex language descriptions. It struggles with spatial descriptions that contain hidden object references and objects described with specific textual annotations, which can affect navigation accuracy. Another limitation lies in the reliance on manually defined thresholds for various navigation conditions like the lightning condition. These thresholds may not generalize well across different environments and scenarios, requiring further investigation into adaptive or learning-based approaches to threshold tuning.
VI-B Conclusion
In this paper, we introduced VL-Nav, a vision-language navigation framework that operates efficiently in real time on resource-constrained platforms. By combining pixel-wise vision-language features with a heuristic exploration strategy, our HVL spatial reasoning method demonstrates robust performance in a variety of indoor and outdoor settings. Empirically, VL-Nav not only achieves real-time navigation at 30 Hz on a Jetson Orin NX but also surpasses existing approaches by 44.15%, achieving an overall success rate of 86.3%. The key to these gains lies in effectively leveraging semantic cues from vision-language embeddings to prioritize frontiers and potential object instances, thereby aligning navigation decisions more closely with human-like reasoning.
Future work will explore extending VL-Nav to handle more complex instructions that involve multi-step tasks, temporal reasoning (e.g., tracking a moving target), and dynamic environments. Additionally, further integration with large language models could enable more nuanced command parsing and open-vocabulary object detection for even broader applicability. We believe these directions will help advance the capability of VLN systems and bring us closer to deploying versatile, robust robotic assistants in real-world settings.
Acknowledgments
This work was partially supported by the Sony Faculty Innovation Award and the DARPA grant DARPA-PS-23-13. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of DARPA.